OpenStudy (dmezzullo):

Solve: negative 1 over 2 x + 3 = −x + 7 1 over 2 8 −8 2 @sue101 @jazy

OpenStudy (anonymous):

\[1/2x + 3 = -x + 7\]

OpenStudy (mayankdevnani):

\[\frac{1}{2x+3}=-x+7\]

OpenStudy (dmezzullo):

?

OpenStudy (anonymous):

wait which way is it supposed to be danny?

OpenStudy (anonymous):

Ummm, is is: \[\frac{ -1 }{ 2x } + 3 = -x + 7\]or\[\frac{ -1 }{ 2x+ 3 } = -x + 7\]

OpenStudy (dmezzullo):